Mr.Chase is 6 years older than 4 times his daughters age. in 7 years, he will be triple her age. How old will Mr.Chase be in 9 years?
Answer by ankor@dixie-net.com(12695)

Write an equation for each statement:
:
Mr.Chase is 6 years older than 4 times his daughters age.
c = 4d + 6
:
in 7 years, he will be triple her age.
c + 7 = 3(d+7)
c + 7 = 3d + 21
c = 3d + 21 - 7
c = 3d + 14
:
Substitute (4d+6) for c and find d:
4d + 6 = 3d + 14
4d - 3d = 14 - 6
d = 8 is the daughters age
:
Find c using the 1st equation
c = 4(8) + 6
c = 38 is Mr Chase's age
:
How old will Mr.Chase be in 9 years?
38 + 9 = 47 in 9 yrs
;
:
Check solutions using the statement,"in 7 years, he will be triple her age."
(38 + 7) = 3(8 + 7)